# Learning Where to Learn: Training Distribution Selection for Provable OOD Performance

### v1

@@ Appendix C Details for section 4Practical algorithms for training distribution design
This appendix expands on the content of Sec. 4 and provides the deferred proofs of the theoretical results from that section. SM C.1 concerns the bilevel minimization algorithm from Sec. 4.1 and SM C.2 pertains to the alternating minimization algorithm from Sec. 4.2.
C.1 Derivations for subsection 4.1Exact bilevel formulation
@@
$\mathcal{G}\mapsto\Psi(\mathcal{G})\coloneqq\frac{1}{2}\lVert\iota_{\nu}\mathcal{G}-\mathcal{G}^{\star}\rVert_{L^{2}_{\nu}(\mathcal{U})}^{2}=\operatorname{\mathbb{E}}_{u\sim\nu}\lvert\mathcal{G}(u)-\mathcal{G}^{\star}(u)\rvert^{2}\,.$ (C.1)
This is the inner objective in (4.1). In the absence of RKHS norm penalization, typically one must assume the existence of minimizers of $\Psi$ [46, Assump. 1, p. 3]. The next lemma characterizes them.
Lemma C.1 (inner minimization solution).
Fix a set $\mathsf{P}\subseteq\mathscr{P}_{2}(\mathcal{U})$ of probability measures. Suppose that for every $\nu\in\mathsf{P}$, there exists $\widehat{\mathcal{G}}^{(\nu)}\in\mathcal{H}$ such that $\Psi(\widehat{\mathcal{G}}^{(\nu)})=\inf_{\mathcal{G}\in\mathcal{H}}\Psi(\mathcal{G})$, where $\Psi$ is as in (C.1). Then $\widehat{\mathcal{G}}^{(\nu)}$ satisfies the compact operator equation $\mathcal{K}_{\nu}\widehat{\mathcal{G}}^{(\nu)}=\iota_{\nu}^{*}\mathcal{G}^{\star}$ in $\mathcal{H}$.
@@
Next, recall the Lagrangian $\mathcal{L}$ from (4.2). We prove Lem. 4.1 on the adjoint state equation.
Proof of Lem. 4.1.
@@
We are now in a position to prove Prop. 4.2, which computes the first variation of $\mathcal{J}\colon\nu\mapsto\mathcal{L}(\nu,\widehat{\mathcal{G}}^{(\nu)},\lambda^{(\nu)}(\widehat{\mathcal{G}}^{(\nu)}))$ in the sense of [48, Def. 7.12, p. 262].
Proof of Prop. 4.2.
The first variation [48, Def. 7.12, p. 262] of $\mathcal{L}$ with respect to the variable $\nu$ is given by the function
@@
that maps from $\mathcal{U}$ to $\mathbb{R}$, which is independent of $\nu$ given $\mathcal{G}$ and $\lambda$. Now let $\widehat{\mathcal{G}}^{(\nu)}$ be as in Lem. C.1 and $\lambda^{(\nu)}$ be as in Lem. 4.1. Due to the smoothness of $\mathcal{L}$ (4.2) with respect to both $\mathcal{G}$ and $\nu$—in particular, $\mathcal{L}$ is convex in $\mathcal{G}$ and linear in $\nu$—the envelope theorem [3] ensures that
@@
Remark C.2 (Wasserstein gradient).
As discussed in Rmk. 4.3, it is clear that
$\nabla D\mathcal{J}=\bigl{(}\nabla\widehat{\mathcal{G}}^{(\nu)}-\nabla\mathcal{G}^{\star}\bigr{)}\lambda^{(\nu)}(\widehat{\mathcal{G}}^{(\nu)})+\bigl{(}\widehat{\mathcal{G}}^{(\nu)}-\mathcal{G}^{\star}\bigr{)}\nabla\bigl{(}\lambda^{(\nu)}(\widehat{\mathcal{G}}^{(\nu)})\bigr{)}$
is the Wasserstein gradient of $\mathcal{J}$ [48, Chp. 8.2]. Its computation requires the usual $\mathcal{U}$-gradient of the trained model $\widehat{\mathcal{G}}^{(\nu)}$, the true map $\mathcal{G}^{\star}$, and the optimal adjoint state $\lambda^{(\nu)}(\widehat{\mathcal{G}}^{(\nu)})$.
We now turn to the parametric training distribution setting. Let $\nu=\nu_{\vartheta}$, where $\vartheta\in\mathcal{V}$ and $(\mathcal{V},\langle\cdot,\cdot\rangle_{\mathcal{V}},\lVert{\,\cdot\,}\rVert_{\mathcal{V}})$ is a parameter Hilbert space as introduced in Sec. 4.1. The main result of Sec. 4.1 is Thm. 4.4. Before proving that theorem, we require the following lemma.
Lemma C.3 (differentiation under the integral).
@@
$\nabla_{\vartheta}\int_{\mathcal{U}}\phi(u)\nu_{\vartheta}(du)=\int_{\mathcal{U}}\phi(u)\Bigl{(}\nabla_{\vartheta}\log p_{\vartheta}(u)\Bigr{)}\nu_{\vartheta}(du)\,.$ (C.2)
@@
With the preceding lemma in hand, we are now able to prove Thm. 4.4.
Proof of Thm. 4.4.
We use the chain rule for Frechét derivatives to compute $D\mathsf{J}\colon\mathcal{V}\to\mathcal{V}^{*}$ and hence the gradient $\nabla\mathsf{J}\colon\mathcal{V}\to\mathcal{V}$. Define $Z\colon\mathcal{V}\to\mathscr{P}_{2}(\mathcal{U})$ by $Z(\vartheta)=\nu_{\vartheta}$. Then Lem. C.3 and duality allow us to identify $DZ$ with the (possibly signed) $\mathcal{V}$-valued measure
@@
Lemma C.4 (model update).
@@
$\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}\coloneqq\sum_{n=1}^{N}\kappa({\,\cdot\,},u_{n})\beta_{n}\approx\widehat{\mathcal{G}}^{(\nu)}\,,\quad\text{where}\quad\beta=\bigl{(}\kappa(U,U)+N\sigma^{2}I_{N}\bigr{)}^{-1}\mathcal{G}^{\star}(U)\in\mathbb{R}^{N}\,.$ (C.3)
@@
$\mathcal{K}_{\nu^{(N)}}h=\frac{1}{N}\sum_{n=1}^{N}\kappa({\,\cdot\,},u_{n})h(u_{n})\eqqcolon\frac{1}{N}\kappa({\,\cdot\,},U)h(U)$ (C.4)
@@
In Lem. C.4, we expect that $\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}\to\widehat{\mathcal{G}}^{(\nu)}$ as $N\to\infty$ and $\sigma\to 0$. This lemma takes care of the finite data approximation of the model. Next, we tackle the adjoints.
@@
Lemma C.5 (adjoint state update).
For $J\in\mathbb{N}$ and a sequence $\{M_{j}\}\subseteq\mathbb{N}$, define $\mathbb{Q}^{(J)}\coloneqq\frac{1}{J}\sum_{j=1}^{J}\delta_{\mu_{j}^{(M_{j})}}$, where $\mu_{j}^{(M_{j})}\coloneqq\frac{1}{M_{j}}\sum_{m=1}^{M_{j}}\delta_{v_{m}^{j}}$, $V^{j}\coloneqq\{v_{m}^{j}\}_{m=1}^{M_{j}}\sim\mu_{j}^{\otimes M_{j}}$, and $\{\mu_{j}\}_{j=1}^{J}\sim\mathbb{Q}^{\otimes J}$. Then with $\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}$ and $U$ as in (C.3), the adjoint values $(\lambda^{(\nu)}(\widehat{\mathcal{G}}^{(\nu)}))(U)\in\mathbb{R}^{N}$ are approximated by
$\widehat{\lambda}_{N,J}^{(\nu,\sigma)}(U)\coloneqq\Bigl{(}\kappa(U,U)+N\sigma^{2}I_{N}\Bigr{)}^{-1}\frac{N}{J}\sum_{j=1}^{J}\frac{1}{M_{j}}\kappa(U,V^{j})\Bigl{(}\mathcal{G}^{\star}(V^{j})-\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}(V^{j})\Bigr{)}\,.$ (C.5)
@@
Let $\mu\sim\mathbb{Q}$, where $\mathbb{Q}\in\mathscr{P}_{2}(\mathscr{P}_{2}(\mathcal{U}))$ is fixed. For some $M$, let $\mu^{(M)}=\frac{1}{M}\sum_{m=1}^{M}\delta_{v_{m}}$ be the empirical measure with $V=\{v_{m}\}_{m=1}^{M}\sim\mu^{\otimes M}$. Consider the case $J=1$ so that $\mathbb{Q}^{(1)}=\delta_{\mu^{(M)}}$; the general case for $J>1$ as asserted in the lemma follows by linearity. With $\nu=\nu^{(N)}$ being the empirical data measure corresponding to $U\in\mathcal{U}^{N}$, $\mathcal{G}=\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}$ being the trained model from Lem. C.4, and $\mathbb{Q}^{(1)}$ in place of $\mathbb{Q}$, the adjoint equation (4.3) from Lem. 4.1 is
@@
By evaluating the preceding display at the $N$ data points $U=\{u_{n}\}_{n=1}^{N}$ and using (C.4), we obtain
@@
This is a linear equation for the unknown $\lambda(U)\in\mathbb{R}^{N}$. Finally, we regularize the equation by replacing $\kappa(U,U)$ with $\kappa(U,U)+\sigma^{2}I_{N}$. We denote the solution of the resulting system by $\widehat{\lambda}_{N,1}^{(\nu,\sigma)}(U)$, which is our desired approximation (C.5) of $(\lambda^{(\nu)}(\widehat{\mathcal{G}}^{(\nu)}))(U)\in\mathbb{R}^{N}$. ∎
In Lem. C.5, the vector $\widehat{\lambda}_{N,J}^{(\nu,\sigma)}(U)\in\mathbb{R}^{N}$ depends on $\nu$ (through the empirical measure $\nu^{(N)}\coloneqq\frac{1}{N}\sum_{n=1}^{N}\delta_{u_{n}}$), the regularization parameter $\sigma$, the current model $\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}$, and the random empirical measure $\mathbb{Q}^{(J)}$ (which itself depends on the sequence $\{M_{j}\}_{j=1}^{J}$). The formula (C.5) requires the inversion of the same regularized kernel matrix as in the model update (C.3). Thus, it is advantageous to pre-compute a factorization of this matrix so that it may be reused in the adjoint update step.
Finally, we perform the numerically implementable training data distribution update by inserting the approximations $\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}$ and $\widehat{\lambda}_{N,J}^{(\nu,\sigma)}(U)$ from Lems. C.4C.5 and empirical measure $\nu_{\vartheta}^{(N)}$ into the exact gradient formula (4.4) for $\nabla\mathsf{J}$. This yields the approximate gradient
$\begin{split}\widehat{\nabla\mathsf{J}}(\vartheta)&\coloneqq\int_{\mathcal{U}}\bigl{(}\widehat{\mathcal{G}}_{N}^{(\nu_{\vartheta},\sigma)}(u)-\mathcal{G}^{\star}(u)\bigr{)}\widehat{\lambda}_{N,J}^{(\nu_{\vartheta},\sigma)}(u)\bigl{(}\nabla_{\vartheta}\log p_{\vartheta}(u)\bigr{)}\nu_{\vartheta}^{(N)}(du)\,,\end{split}$ (C.6)
where $\nu_{\vartheta}^{(N)}$ is the empirical measure of the same training data points used to obtain both the trained model and adjoint state vector, and we identify the vector $\widehat{\lambda}_{N,J}^{(\nu,\sigma)}(U)\in\mathbb{R}^{N}$ with a function $\widehat{\lambda}_{N,J}^{(\nu,\sigma)}\in L^{2}_{\nu_{\vartheta}^{(N)}}$. This leads to Alg. 3. In this practical algorithm, we allow for a varying number $N_{k}$ of training samples from $\nu_{\vartheta^{(k)}}$ and a varying regularization strength $\sigma_{k}$ at each iteration $k$ in Line 4. The integrals in (C.6) and Line 4 of Alg. 3 are with respect to the empirical measure; hence, they equal the equally-weighted average over the finite data samples that make up the empirical measure.
@@
$\vartheta^{(k+1)}=\vartheta^{(k)}-t_{k}\int_{\mathcal{U}}\Bigl{(}\widehat{\mathcal{G}}_{N_{k}}^{(\nu_{\vartheta^{(k)}},\sigma_{k})}(u)-\mathcal{G}^{\star}(u)\Bigr{)}\widehat{\lambda}_{N_{k},J}^{(\nu_{\vartheta^{(k)}},\sigma_{k})}(u)\bigl{(}\nabla_{\vartheta}\log p_{\vartheta^{(k)}}(u)\bigr{)}\nu_{\vartheta^{(k)}}^{(N_{k})}(du)$

### v3

@@ Appendix D Details for section 4Practical algorithms for training distribution design
This appendix expands on the content of Sec. 4 and provides the deferred proofs of the theoretical results from that section. It begins in SM D.1 by identifying a degeneracy of the bilevel problem formulation when infinite data are available. SM D.2 concerns the bilevel minimization algorithm from Sec. 4.1 and SM D.3 pertains to the alternating minimization algorithm from Sec. 4.2.
D.1 Necessity of finite data
Although Secs. 3 and 4 are formulated at the infinite data level, i.e., with full access to $\mathbb{Q}$ and the candidate training distribution $\nu$, the proposed formulation is only advantageous when $\nu$ is accessible from a finite number of samples. Of course, this is the case in practice (and moreover $\mathbb{Q}$ will often only be accessible from samples as well); recall Rem. 4.1. Our optimize-then-discretize approach of Secs. 3 and 4 simplifies the exposition and notation, while allowing us to derive continuum algorithms that still work at the finite data level. Regardless, it is instructive to see how the necessity of finite data emerges from a theoretical viewpoint.
To this end, recall the functional $\mathcal{E}$$_{\mathbb{Q}}$ from (3.3). Define the infinite mixture probability measure
$\nu_{\mathbb{Q}}\coloneqq\mathbb{E}_{\nu^{\prime}\sim\mathbb{Q}}[\nu^{\prime}]$ (D.1)
via duality with bounded continuous test functions. By linearity, one can check that
$\mathcal{E}_{\mathbb{Q}}(\mathcal{G})=\mathbb{E}_{u^{\prime}\sim\nu_{\mathbb{Q}}}\lVert\mathcal{G}^{\star}(u^{\prime})-\mathcal{G}(u^{\prime})\rVert_{\mathcal{Y}}^{2}\,.$ (D.2)
This display shows that the average OOD accuracy with respect to $\mathbb{Q}$ is equal to the standard test error with respect to the test mixture distribution $\nu_{\mathbb{Q}}$.
Now define the optimal value $V^{\star}_{\mathbb{Q}}\coloneqq\inf_{\mathcal{G}\in\mathcal{H}}\mathcal{E}_{\mathbb{Q}}(\mathcal{G})$. By the definition of the trained model in (4.1), it holds that $V^{\star}_{\mathbb{Q}}=\mathcal{E}_{\mathbb{Q}}(\widehat{\mathcal{G}}^{(\nu_{\mathbb{Q}})})$. The inequality $V^{\star}_{\mathbb{Q}}\leq\inf_{\nu}\mathcal{E}_{\mathbb{Q}}(\widehat{\mathcal{G}}^{(\nu)})$ holds because $V^{\star}_{\mathbb{Q}}$ is the optimal value over all elements from $\mathcal{H}$. Moreover, $\inf_{\nu}\mathcal{E}_{\mathbb{Q}}(\widehat{\mathcal{G}}^{(\nu)})\leq\mathcal{E}_{\mathbb{Q}}(\widehat{\mathcal{G}}^{(\nu_{\mathbb{Q}})})=V^{\star}_{\mathbb{Q}}$ due to optimality over training distributions. We deduce that
$\inf_{\nu}\mathcal{E}_{\mathbb{Q}}(\widehat{\mathcal{G}}^{(\nu)})=\inf_{\mathcal{G}\in\mathcal{H}}\mathcal{E}_{\mathbb{Q}}(\mathcal{G})\,.$ (D.3)
That is, the optimal value of the objective over distributions versus the objective over models coincide. Write $\widehat{\nu}$ for a solution of $\inf_{\nu}\mathcal{E}_{\mathbb{Q}}(\widehat{\mathcal{G}}^{(\nu)})$ from (4.1). The preceding display then says
$\mathcal{E}_{\mathbb{Q}}(\widehat{\mathcal{G}}^{(\widehat{\nu})})=\mathcal{E}_{\mathbb{Q}}(\widehat{\mathcal{G}}^{(\nu_{\mathbb{Q}})})\,.$ (D.4)
This does not imply that $\widehat{\nu}$ equals $\nu_{\mathbb{Q}}$. Indeed, these infinite-dimensional optimization problems over possibly noncompact sets could admit multiple minima. What (D.4) does suggest is that the mixture $\nu_{\mathbb{Q}}$ (which only depends on $\mathbb{Q}$) is an equally good solution for making the OOD accuracy small compared to the adaptive measure $\widehat{\nu}$ (which depends on $\mathcal{G}^{\star}$, $\mathcal{H}$, and $\mathbb{Q}$), assuming full knowledge of $\mathbb{Q}$ and infinite training data. In this case, there is no reason to construct $\widehat{\nu}$ if $\nu_{\mathbb{Q}}$ is known.
However, for each $N$, writing $\nu^{N}=\frac{1}{N}\sum_{n=1}^{N}\delta_{u_{n}}$ for the empirical measure corresponding to i.i.d. samples $u_{n}\sim\nu$, we can instead consider
$\widehat{\nu}_{N}\in\operatorname*{arg\,min}{\mathopen{}\mathclose{{\left\{\mathscr{E}_{\mathbb{Q}}\bigl(\widehat{\mathcal{G}}_{N}^{(\nu)}\bigr)\,\middle|\,\nu\in\mathsf{P}\subseteq\mathscr{P}_{2}(\mathcal{U})\quad\text{and}\quad\widehat{\mathcal{G}}_{N}^{(\nu)}\in\operatorname*{arg\,min}_{\mathcal{G}\in\mathcal{H}}\operatorname{\mathbb{E}}_{u\sim\nu^{N}}\big\lVert\mathcal{G}^{\star}(u)-\mathcal{G}(u)\big\rVert_{\mathcal{Y}}^{2}}}\right\}}$ (D.5)
for some search space $\mathsf{P}$. The preceding argument leading to the degenerate behavior shown in (D.4) no longer holds for the empirical bilevel problem (D.5) due to the finite data constraint. It is now possible that models trained on samples from $\widehat{\nu}_{N}$ can outperform models trained on samples from the mixture $\nu_{\mathbb{Q}}$. Indeed, we observe the superiority of $\widehat{\nu}_{N}$ over $\nu_{\mathbb{Q}}$ even for simple Gaussian parametrizations in Secs. 5 and F.1. More generally, the test distribution $\nu_{\mathbb{Q}}$ is not the optimal distribution from which to generate training data for regression problems. This fact is known to the numerical analysis community [1, 21], building off of variance reduction and importance sampling ideas [5] for Monte Carlo and quasi-Monte Carlo methods [15, 27]. It remains true for (linear) operator learning, where de Hoop et al. [24] suggests that one should pick a training data distribution over input functions whose samples are rougher than function samples from the test distribution. This implies that the training distribution support should (strictly) contain the support of the test distribution. However, there are fundamental limits on how much improvement one can expect to obtain by adjusting the training data distribution [42, 53].
Another practical takeaway from the preceding discussions is that if data from $\mathbb{Q}$ are very limited and fixed once and for all, then a finite sample approximation will not capture the mixture $\nu_{\mathbb{Q}}$ well. This could lead to a poor performing model when trained on the aggregate empirical mixture samples. On the other hand, the proposed approach delivers a distribution $\widehat{\nu}_{N}$ that ideally can be sampled from at will (leading to many more training samples, e.g., a generative model) and uses information about $\mathcal{G}^{\star}$ and $\mathcal{H}$ to compensate for limited information about $\mathbb{Q}$.
D.2 Derivations for subsection 4.1Exact bilevel formulation
@@ D.1 Necessity of finite data
$\mathcal{G}\mapsto\Psi(\mathcal{G})\coloneqq\frac{1}{2}\lVert\iota_{\nu}\mathcal{G}-\mathcal{G}^{\star}\rVert_{L^{2}_{\nu}(\mathcal{U})}^{2}=\operatorname{\mathbb{E}}_{u\sim\nu}\lvert\mathcal{G}(u)-\mathcal{G}^{\star}(u)\rvert^{2}\,.$ (D.6)
This is the inner objective in (4.1). In the absence of RKHS norm penalization, typically one must assume the existence of minimizers of $\Psi$ [79, Assump. 1, p. 3]. The next lemma characterizes them.
Lemma D.1 (inner minimization solution).
Fix a set $\mathsf{P}\subseteq\mathscr{P}_{2}(\mathcal{U})$ of probability measures. Suppose that for every $\nu\in\mathsf{P}$, there exists $\widehat{\mathcal{G}}^{(\nu)}\in\mathcal{H}$ such that $\Psi(\widehat{\mathcal{G}}^{(\nu)})=\inf_{\mathcal{G}\in\mathcal{H}}\Psi(\mathcal{G})$, where $\Psi$ is as in (D.6). Then $\widehat{\mathcal{G}}^{(\nu)}$ satisfies the compact operator equation $\mathcal{K}_{\nu}\widehat{\mathcal{G}}^{(\nu)}=\iota_{\nu}^{*}\mathcal{G}^{\star}$ in $\mathcal{H}$.
@@
Next, recall the Lagrangian $\mathcal{L}$ from (4.2). We prove Lem. 4.2 on the adjoint state equation.
Proof of Lem. 4.2.
@@
We are now in a position to prove Prop. 4.3, which computes the first variation of $\mathcal{J}\colon\nu\mapsto\mathcal{L}(\nu,\widehat{\mathcal{G}}^{(\nu)},\lambda^{(\nu)}(\widehat{\mathcal{G}}^{(\nu)}))$ in the sense of Santambrogio [82, Def. 7.12, p. 262].
Proof of Prop. 4.3.
The first variation [82, Def. 7.12, p. 262] of $\mathcal{L}$ with respect to the variable $\nu$ is given by the function
@@
that maps from $\mathcal{U}$ to $\mathbb{R}$, which is independent of $\nu$ given $\mathcal{G}$ and $\lambda$. Now let $\widehat{\mathcal{G}}^{(\nu)}$ be as in Lem. D.1 and $\lambda^{(\nu)}$ be as in Lem. 4.2. Due to the smoothness of $\mathcal{L}$ (4.2) with respect to both $\mathcal{G}$ and $\nu$—in particular, $\mathcal{L}$ is convex in $\mathcal{G}$ and linear in $\nu$—the envelope theorem [4] ensures that
@@
Remark D.2 (Wasserstein gradient).
As discussed in Rmk. 4.4, it is clear that
$\nabla\bigl[(D\mathcal{J})(\nu)\bigr]=\bigl(\nabla\widehat{\mathcal{G}}^{(\nu)}-\nabla\mathcal{G}^{\star}\bigr)\lambda^{(\nu)}(\widehat{\mathcal{G}}^{(\nu)})+\bigl(\widehat{\mathcal{G}}^{(\nu)}-\mathcal{G}^{\star}\bigr)\nabla\bigl(\lambda^{(\nu)}(\widehat{\mathcal{G}}^{(\nu)})\bigr)$
is the Wasserstein gradient of $\mathcal{J}$ [82, Chp. 8.2]. Its computation requires the usual $\mathcal{U}$-gradient of the trained model $\widehat{\mathcal{G}}^{(\nu)}$, the true map $\mathcal{G}^{\star}$, and the optimal adjoint state $\lambda^{(\nu)}(\widehat{\mathcal{G}}^{(\nu)})$.
We now turn to the parametric training distribution setting. Let $\nu=\nu_{\vartheta}$, where $\vartheta\in\mathcal{V}$ and $(\mathcal{V},\langle\cdot,\cdot\rangle_{\mathcal{V}},\lVert{\,\cdot\,}\rVert_{\mathcal{V}})$ is a parameter Hilbert space as introduced in Sec. 4.1. The main result of Sec. 4.1 is Thm. 4.5. Before proving that theorem, we require the following lemma.
Lemma D.3 (differentiation under the integral).
@@
$\nabla_{\vartheta}\int_{\mathcal{U}}\phi(u)\nu_{\vartheta}(du)=\int_{\mathcal{U}}\phi(u)\Bigl(\nabla_{\vartheta}\log p_{\vartheta}(u)\Bigr)\nu_{\vartheta}(du)\,.$ (D.7)
@@
With the preceding lemma in hand, we are now able to prove Thm. 4.5.
Proof of Thm. 4.5.
We use the chain rule for Frechét derivatives to compute $D\mathsf{J}\colon\mathcal{V}\to\mathcal{V}^{*}$ and hence the gradient $\nabla\mathsf{J}\colon\mathcal{V}\to\mathcal{V}$. Define $Z\colon\mathcal{V}\to\mathscr{P}_{2}(\mathcal{U})$ by $Z(\vartheta)=\nu_{\vartheta}$. Then Lem. D.3 and duality allow us to identify $DZ$ with the (possibly signed) $\mathcal{V}$-valued measure
@@
Lemma D.4 (model update).
@@
$\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}\coloneqq\sum_{n=1}^{N}\kappa({\,\cdot\,},u_{n})\beta_{n}\approx\widehat{\mathcal{G}}^{(\nu)}\,,\quad\text{where}\quad\beta=\bigl(\kappa(U,U)+N\sigma^{2}I_{N}\bigr)^{-1}\mathcal{G}^{\star}(U)\in\mathbb{R}^{N}\,.$ (D.8)
@@
$\mathcal{K}_{\nu^{(N)}}h=\frac{1}{N}\sum_{n=1}^{N}\kappa({\,\cdot\,},u_{n})h(u_{n})\eqqcolon\frac{1}{N}\kappa({\,\cdot\,},U)h(U)$ (D.9)
@@
In Lem. D.4, we expect that $\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}\to\widehat{\mathcal{G}}^{(\nu)}$ as $N\to\infty$ and $\sigma\to 0$. This lemma takes care of the finite data approximation of the model. Next, we tackle the adjoints.
@@
Lemma D.5 (adjoint state update).
For $J\in\mathbb{N}$ and a sequence $\{M_{j}\}\subseteq\mathbb{N}$, define $\mathbb{Q}^{(J)}\coloneqq\frac{1}{J}\sum_{j=1}^{J}\delta_{\mu_{j}^{(M_{j})}}$, where $\mu_{j}^{(M_{j})}\coloneqq\frac{1}{M_{j}}\sum_{m=1}^{M_{j}}\delta_{v_{m}^{j}}$, $V^{j}\coloneqq\{v_{m}^{j}\}_{m=1}^{M_{j}}\sim\mu_{j}^{\otimes M_{j}}$, and $\{\mu_{j}\}_{j=1}^{J}\sim\mathbb{Q}^{\otimes J}$. Then with $\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}$ and $U$ as in (D.8), the adjoint values $(\lambda^{(\nu)}(\widehat{\mathcal{G}}^{(\nu)}))(U)\in\mathbb{R}^{N}$ are approximated by
$\widehat{\lambda}_{N,J}^{(\nu,\sigma)}(U)\coloneqq\Bigl(\kappa(U,U)+N\sigma^{2}I_{N}\Bigr)^{-1}\frac{N}{J}\sum_{j=1}^{J}\frac{1}{M_{j}}\kappa(U,V^{j})\Bigl(\mathcal{G}^{\star}(V^{j})-\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}(V^{j})\Bigr)\,.$ (D.10)
@@
Let $\mu\sim\mathbb{Q}$, where $\mathbb{Q}\in\mathscr{P}_{2}(\mathscr{P}_{2}(\mathcal{U}))$ is fixed. For some $M$, let $\mu^{(M)}=\frac{1}{M}\sum_{m=1}^{M}\delta_{v_{m}}$ be the empirical measure with $V=\{v_{m}\}_{m=1}^{M}\sim\mu^{\otimes M}$. Consider the case $J=1$ so that $\mathbb{Q}^{(1)}=\delta_{\mu^{(M)}}$; the general case for $J>1$ as asserted in the lemma follows by linearity. With $\nu=\nu^{(N)}$ being the empirical data measure corresponding to $U\in\mathcal{U}^{N}$, $\mathcal{G}=\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}$ being the trained model from Lem. D.4, and $\mathbb{Q}^{(1)}$ in place of $\mathbb{Q}$, the adjoint equation (4.3) from Lem. 4.2 is
@@
By evaluating the preceding display at the $N$ data points $U=\{u_{n}\}_{n=1}^{N}$ and using (D.9), we obtain
@@
This is a linear equation for the unknown $\lambda(U)\in\mathbb{R}^{N}$. Finally, we regularize the equation by replacing $\kappa(U,U)$ with $\kappa(U,U)+\sigma^{2}I_{N}$. We denote the solution of the resulting system by $\widehat{\lambda}_{N,1}^{(\nu,\sigma)}(U)$, which is our desired approximation (D.10) of $(\lambda^{(\nu)}(\widehat{\mathcal{G}}^{(\nu)}))(U)\in\mathbb{R}^{N}$. ∎
In Lem. D.5, the vector $\widehat{\lambda}_{N,J}^{(\nu,\sigma)}(U)\in\mathbb{R}^{N}$ depends on $\nu$ (through the empirical measure $\nu^{(N)}\coloneqq\frac{1}{N}\sum_{n=1}^{N}\delta_{u_{n}}$), the regularization parameter $\sigma$, the current model $\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}$, and the random empirical measure $\mathbb{Q}^{(J)}$ (which itself depends on the sequence $\{M_{j}\}_{j=1}^{J}$). The formula (D.10) requires the inversion of the same regularized kernel matrix as in the model update (D.8). Thus, it is advantageous to pre-compute a factorization of this matrix so that it may be reused in the adjoint update step.
Finally, we perform the numerically implementable training data distribution update by inserting the approximations $\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}$ and $\widehat{\lambda}_{N,J}^{(\nu,\sigma)}(U)$ from Lems. D.4D.5 and empirical measure $\nu_{\vartheta}^{(N)}$ into the exact gradient formula (4.4) for $\nabla\mathsf{J}$. This yields the approximate gradient
[…]
where $\nu_{\vartheta}^{(N)}$ is the empirical measure of the same training data points used to obtain both the trained model and adjoint state vector, and we identify the vector $\widehat{\lambda}_{N,J}^{(\nu,\sigma)}(U)\in\mathbb{R}^{N}$ with a function $\widehat{\lambda}_{N,J}^{(\nu,\sigma)}\in L^{2}_{\nu_{\vartheta}^{(N)}}$. This leads to Alg. 3. In this practical algorithm, we allow for a varying number $N_{k}$ of training samples from $\nu_{\vartheta^{(k)}}$ and a varying regularization strength $\sigma_{k}$ at each iteration $k$ in Line 4. The integrals in (D.11) and Line 4 of Alg. 3 are with respect to the empirical measure; hence, they equal the equally-weighted average over the finite data samples that make up the empirical measure.
@@
$\vartheta^{(k+1)}=\vartheta^{(k)}-t_{k}\int_{\mathcal{U}}\Bigl(\widehat{\mathcal{G}}_{N_{k}}^{(\nu_{\vartheta^{(k)}},\sigma_{k})}(u)-\mathcal{G}^{\star}(u)\Bigr)\widehat{\lambda}_{N_{k},J}^{(\nu_{\vartheta^{(k)}},\sigma_{k})}(u)\bigl(\nabla_{\vartheta}\log p_{\vartheta^{(k)}}(u)\bigr)\nu_{\vartheta^{(k)}}^{(N_{k})}(du)\$